I am supposed to complete the square: x^2 + 5/2x = 5
How do I get rid of the fraction 5/2- I can't square it, can I? I don't know how to solve it.

Answer by stanbon(75887)

complete the square: x^2 + 5/2x = 5
Multiply the 5/2 by 1/2; square the result and add to both sides to get:
=============================
x^2 + (5/2)x + (5/4)^2 = 5 + (5/4)^2
---
(x+(5/4))^2 = (80/16) + (25/16) = 105/16
---
Take the square root of both sides to get:
x+(5/4) = +/-(sqrt(105))/4
---
x = [-5+/-(sqrt(105)]/4
===================================
Cheers,
Stan H.